OpenStudy (anonymous):

Solve. x2 + 8x + 25 = 0 please show work...ty

OpenStudy (johnweldon1993):

if you aren't very good at factoring......have you heard of the quadratic equation?

OpenStudy (johnweldon1993):

\[\frac{ -b \pm \sqrt{b² - 4ac} }{ 2a }\] where a = coefficient of first term = 1 b = coefficient of second term = 8 c = third term = 25 plug those into that equation and simplify hint****if you aren't dealing with complex numbers yet....then there is no answer √8² - 4(1)(25) √64 - 100 √of a negative number is...a complex number again...if you arent dealing with these yet...then there is no solution

OpenStudy (anonymous):

i got –4 ± i

OpenStudy (anonymous):

Use the quadratic formula, where: a = 1, b = 8, c = 25 -8 +- sqrt(8^2 - 4(1)(25))/2(1) -8 +- sqrt(64 - 100)/2 -8 +- sqrt(-36)/2 Now we have a negative in our square root, so factor out an imaginary. -8 +- i*sqrt(36)/2 Continuing: -8 +- 6i/2 -4 +- 3i So your answer is -4 +- 3i

OpenStudy (anonymous):

but its wrong...what did I do wrong i forgot to bring down the negative but I missed a whole step...ty ..watavi

OpenStudy (anonymous):

If you could show your work for what you did, I could help you